OpenStudy (anonymous):

will give medal how would i graph -1/2x + 1/6y = 2/3

OpenStudy (anonymous):

and

OpenStudy (tkhunny):

What is your most familiar form of a line? Slope-Intercept?

OpenStudy (anonymous):

@agent0smith

OpenStudy (tkhunny):

You don't have a favorite form with which you are familiar? I've never heard of the agent0smith form.

OpenStudy (anonymous):

no i do not have a favorite form

OpenStudy (agent0smith):

\[\Large -\frac{ 1 }{ 2 } x + \frac{ 1 }{ 6 } y = \frac{ 2 }{3 }\]The easiest method to graph from this form would be to find x- and y- intercepts, or put it into y=mx+b form if you like.

OpenStudy (anonymous):

the first way works

OpenStudy (agent0smith):

\[\Large -\frac{ 1 }{ 2 } x + \frac{ 1 }{ 6 } y = \frac{ 2 }{3 }\] To find the x-intercept: plug in y=0 and solve the equation for x To find the y-intercept: plug in x=0 and solve the equation for y

OpenStudy (shamil98):

the agent0smith form, i lol'd

OpenStudy (shamil98):

Try converting it to slope-intercept form or do what agent0smith suggested. y = mx +b m = slope (rise/run of your line) b = y-intercept (where you start graphing from on the y-axis),

OpenStudy (anonymous):

im confused

OpenStudy (agent0smith):

Start with the x-intercept. To find the x-intercept: plug in y=0 in the following equation, and solve for x \[\Large -\frac{ 1 }{ 2 } x + \frac{ 1 }{ 6 } y = \frac{ 2 }{3 }\]

OpenStudy (anonymous):

-4/3

OpenStudy (anonymous):

?

OpenStudy (agent0smith):

Yes, so your x-int is \[\Large \left( -\frac{ 4 }{ 3}, 0 \right)\] To find the y-intercept: plug in x=0 and solve the equation for y

OpenStudy (anonymous):

y = 4

OpenStudy (agent0smith):

So your y intercept is (0, 4). Now plot this point, and the x-intercept above on a graph.

OpenStudy (agent0smith):

Once you plot the two points, draw a straight line through both, with arrows on each end.

OpenStudy (anonymous):

i don't know how to graph

OpenStudy (tkhunny):

-1/2x + 1/6y = 2/3 Maybe, multiply by 6 -3x + y = 4 Add 3x to Slope-Intercept Form y = 3x + 4 Divide by 4 to Intercept Form \(\dfrac{x}{-4/3} + \dfrac{y}{4} = 1\) You should be familiar with the various "Forms". One or two forms may be most valuable, but under exam time pressure, the quickest and easiest solutions come from knowing the form that matches the given data most closely.

OpenStudy (anonymous):

im not sure which graph i correct

OpenStudy (tkhunny):

That "Intercept Form" is a GREAT hint! Which one has x-intercept x = -4/3 and y-intercept y = 4?

OpenStudy (anonymous):

im not sure

OpenStudy (tkhunny):

You cannot pass this course if you cannot read an x-y coordinate system. Do you live in a community where the city blocks are laid out in a rectangular grid? Generally, these things start at (0,0), known as the origin. x increases in the positive direction to the RIGHT. y increases in the positive direction UP. In both cases, the opposite direction is the mirror image but in negative numbers. (0,4) is 0 from the Origin in the right or left direction and 4 UP (positive y-direction.) (-4/3,) is 4/3 in the negative x-direction and 0 up or down.

OpenStudy (agent0smith):

^Yeah I'm hoping you're able to deal with coordinates. The y-intercept is where the graph touches the y-axis, where the two lines meet (intersect). The x-intercept is where the graph touches the x-axis.

OpenStudy (anonymous):

would it be the first graph

OpenStudy (tkhunny):

Where does the first graph cross the x-axis (the horizontal axis)?

OpenStudy (usukidoll):